OpenStudy (anonymous):

HELP PLEASE? Bayside insurance offers two health plans.Under plan A Giselle would have to pay the first $80 of her medical bills plus 40% of the rest. Under plan B Giselle would pay the first $210, but only 30% of the rest. for what amount of medical bills will plan B save Giselle money? Assume she has over $210 in Bills

OpenStudy (anonymous):

Seems like you aren't the only one having problems getting help today. More people asking questions than answering or helping answer.

OpenStudy (anonymous):

I agree. I have had this problem all day.

OpenStudy (anonymous):

So you need to look at it as an equation: We have M + d(C-M) assuming that C > M, with C = cost, d = percent to pay, M equals minimal cost Does that make sense?

OpenStudy (anonymous):

Yes kind of...That is the formula we use to solve the problem?

OpenStudy (anonymous):

ooooooooh, this is more interesting than that, sorry I misread

OpenStudy (anonymous):

but, you can solve it with just a little more work, that's almost enough

OpenStudy (anonymous):

I have a hard time with word problems, never fails. Especially in algebra

OpenStudy (anonymous):

ahh, I totally understand, it's very common

OpenStudy (anonymous):

so, you need to find when the two cost the same amount of money

OpenStudy (anonymous):

so you have M1 + d1(C-M1) = M2 - d2(C-M2)

OpenStudy (anonymous):

where M1 and d1 are one of the combinations of percent to pay and minimum, M2 and d2 are the other

OpenStudy (anonymous):

if you want help making sense of that formula I can help

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

back, this is a hard problem for many students, but it is doable and understanding the why more than just the answer will be very helpful in later problems

OpenStudy (anonymous):

I think I am way more confused now!

OpenStudy (anonymous):

don't worry, that equation looks WAY more confusing than it is

OpenStudy (anonymous):

I'll fill it in with the numbers, then it will looks less intimidating

OpenStudy (anonymous):

I figured. HAHA

OpenStudy (anonymous):

thank you that might help

OpenStudy (anonymous):

80 + 0.40*(C-80) = 210 + 0.30*(C-210)

OpenStudy (anonymous):

so do I add 80+0.40 first and then 210+0.30 ?

OpenStudy (anonymous):

not quite, that up there is just an algebraic equation, you need to solve for C

OpenStudy (anonymous):

oh boy. LOL

OpenStudy (anonymous):

I'll get rid of the parentheses, 80 + .40*C + 24.0 = 210 + .30C - 63

OpenStudy (anonymous):

I have no clue where to start.

OpenStudy (anonymous):

oh, dang it I screwed up a sign, rather 80 + .40*C - 24.0 = 210 + .30C - 63

OpenStudy (anonymous):

Still have no clue

OpenStudy (anonymous):

So just with the equation, move C to one side like this 80 + .40*C - 24.0 (- .40*C) = 210 + .30*C - 63 (- .40*C) gives you 80 - 24 = 210 -.10*C-63

OpenStudy (anonymous):

then move the rest of the equation to the other side (everything but the C) 80 - 24 + 63 - 210 = 210 - .10*C - 63 (+63 - 210) -91 = -.10*C

OpenStudy (anonymous):

ok I think I see that.